How do you compute the force of attraction between a +1.60 × 10 C charge and a -2.09 × 10C charge if they are 4.01 × 10 m apart?

1
Expert's answer
2022-02-19T04:56:03-0500

Given:

"q_1=1.6*10\:\rm C"

"q_2=-2.9*10\:\rm C"

"r=4.01*10\:\rm m"


The Coulomb's law says



"F=k\frac{|q_1||q_2|}{r^2}""F=9*10^9\frac{16*29}{40.1^2}=2.6*10^9 \:\rm N"

The force is attraction.